# Measuring Psychological Depth in Language Models

### v1

@@ Abstract
Evaluations of creative stories generated by large language models (LLMs) often focus on objective properties of the text, such as its style, coherence, and toxicity. While these metrics are indispensable, they do not speak to a story’s subjective, psychological impact from a reader’s perspective. We introduce the Psychological Depth Scale (PDS), a novel framework rooted in literary theory that measures an LLM’s ability to produce authentic and narratively complex stories that provoke emotion, empathy, and engagement. We empirically validate our framework by showing that humans can consistently evaluate stories based on PDS ( Krippendorff’s alpha). We also explore techniques for automating the PDS to easily scale future analyses. GPT-4o, combined with a novel Mixture-of-Personas (MoP) prompting strategy, achieves an average Spearman correlation of with human judgment while Llama-3-70B scores as high as for empathy. Finally, we compared the depth of stories authored by both humans and LLMs. Surprisingly, GPT-4 stories either surpassed or were statistically indistinguishable from highly-rated human-written stories sourced from Reddit. By shifting the focus from text to reader, the Psychological Depth Scale is a validated, automated, and systematic means of measuring the capacity of LLMs to connect with humans through the stories they tell.
@@
Fabrice Harel-Canada Hanyu Zhou Sreya Mupalla Zeynep Yildiz Amit Sahai Nanyun Peng University of California, Los Angeles fabricehc@cs.ucla.edu
@@ 1 Introduction
Stories play a crucial role in our understanding of ourselves and the world around us Langer (1942); Kaniss (1991). As large language models (LLMs) are increasingly deployed in narrative design and creation, their growing impact on how stories are told calls for a deeper understanding of their narrative power. Current evaluations of LLM-generated stories often focus on objective properties of the text such as discourse structure Liu et al. (2024), fluency Gómez-Rodríguez and Williams (2023), style Wang et al. (2024b), creativity Chakrabarty et al. (2023), diversity Gehrmann et al. (2021), toxicity and bias Wang et al. (2024a). However, it is crucial to extend these evaluations to accommodate the subjective, psychological impact stories have on readers. While some studies have recently explored aspects like empathy Ma et al. (2020); Sabour et al. (2022) and engagement Ghazarian et al. (2020); Xu et al. (2022), they do not fully capture the multifaceted and interconnected nature of the reading experience.
@@
To empirically validate the PDS, we conducted a study involving non-expert humans, who are increasingly engaging with creative content generated by LLMs. We enlisted five undergraduate students from a US public university to analyze a dataset of 97 stories authored by humans and five contemporary LLMs. The raters provided psychological depth ratings and predictions on human or machine authorship, with detailed justifications. This approach addresses three critical research questions to explore the intersection of LLMs and the psychological nuances of creative writing.
@@
To what extent can psychological depth be measured automatically? Leveraging our novel Mixture-of-Personas prompting strategy, GPT-4o achieved a Spearman correlation of with human judgment while Llama-3-70B attains correlations as high as for empathy and for narrative complexity. These results highlight that while no single LLM excels at predicting all components of psychological depth, a strategic combination of different LLMs shows significant promise for automating PDS analyses.
@@
Overall, our findings validate the Psychological Depth Scale as an effective, automated, and systematic means of measuring how well LLMs connect with humans through storytelling. Remarkably, our results reveal that GPT-4 already matches or exceeds the quality of respected stories from Reddit, with of readers believing GPT-4’s stories to be human-written.111Code and dataset will be released upon acceptance.
@@ 2 The Psychological Depth Scale
The metrics for PDS are derived from an extensive literature review within cognitive psychology, media studies, and narrative analysis. 95 peer-reviewed articles and books were included in the final review. 143 evaluative components were extracted from the included works. We employed thematic analysis to group these candidates under five metrics: emotion provocation, empathy, engagement, authenticity, and narrative complexity. Below we analyze these metrics and discuss their impact. We note that while factors contributing to each metric may be complex, measuring their narrative achievement is relatively simple. As PDS is a reader-centered assessment tool, each metric is such that readers will know when a story achieves it. Our surveys reflect this fact.
Emotion Provocation (PROV) measures the narrative’s ability to elicit strong emotional responses, positive or negative. Theories of emotions range across neuroscience Clark-Polner et al. (2016); Mendes (2016), cognitive psychology Ortony et al. (1988); Lazarus (1991), and philosophy Sartre (1971); Nussbaum (2004). Contemporary approaches describe emotions as mental states marked by valence (positivity or negativity) and arousal (level of intensity) Barrett et al. (2016). Narrative evocation of emotions foster attention and interest, and evoked emotions may persist or recur after reading Mar et al. (2011).
Recent fMRI research shows that congruent (i.e. positive valence - high intensity, negative valence - low intensity) textual emotive expressions are more cognitively effective than conflicting (i.e. positive valence - low intensity, negative valence - high intensity) ones Citron et al. (2014); Megalakaki et al. (2019). The disparity between congruent and conflicting narrative emotions may be a factor in "compassion fatigue" Maier (2015); Kinnick et al. (1996) and "psychic numbing" Maier et al. (2016); Slovic (2007) exhibited by readers towards news about mass tragedies. Enhancing a story’s emotional impact is therefore more complex than increasing the amount and intensity of emotion-laden content as this may elicit weaker emotional responses. Given its impact and complexity, we contend that emotional provocation is an achievement of psychologically deep stories and is thereby a crucial metric in narrative quality assessment.
@@
Narrative Complexity (NCOM) refers to the presence of rich and intricate storylines and character development, especially those that engender puzzled intrigue from the reader Somasundaran et al. (2018). Narrative complexity is often achieved through creative techniques such as nonlinear narration, plot twists, and double perspectivation Kiss and Willemsen (2017). Complex narratives present immersive puzzles that motivate readers to undertake mental restructuring and retroactive revision Kiss and Willemsen (2017); Cutting (2019). Exerting cognitive effort facilitates reader attention and interest Empson (1947); Steiner (1978), and can result in a more enjoyable reading experience Zunshine (2006). Narrative complexity is not narrational complexity: simple stories can be obscurely narrated, while complex stories can be free of complex language Cutting (2019). Furthermore, the extent to which narrative complexity is intersubjective depends on situational similarities and contextual cues Tikka and Kaipainen (2017). Thus, narrative complexity is a dynamic and reciprocal measure that goes beyond stylistic and structural choices.
@@ 3.1 Human Stories
We collected human-authored stories from Reddit’s r/WritingPrompts,222https://www.reddit.com/r/WritingPrompts a popular online community with over 18 million users. This forum was chosen for its accessibility, thematic diversity, and the structured nature of its writing prompts. Aspiring writers respond to these prompts (called "premises") with their stories, which readers can upvote or downvote. While these stories may not represent the pinnacle of human writing, Reddit’s voting system allows us to approximate where LLM performance stands relative to variances in human writing quality. We categorized the stories based on their ranking position: Human-Advanced for top-voted stories, Human-Intermediate for medium-voted stories, and Human-Novice for low-voted stories, with average upvotes of 1434, 263, and 9, respectively. To ensure that quality was the primary factor influencing votes, we only selected stories that were posted within 24 hours of one another.
@@ 3.2 LLM Stories
When generating stories with LLMs, we employed a multifaceted approach involving five models, two prompting strategies, and three sampled generations. We intentionally restricted the model architecture to examine the impact of model size on psychological depth. In particular, we chose the Llama-2 family Touvron et al. (2023) with a variety of sizes (7B, 13B, and 70B). We added Vicuna-33B Chiang et al. (2023), which is fine-tuned based on Llama, to fill the size gap between 13B and 70B. We also included GPT-4 OpenAI et al. (2023) as the highest-performing LLM at the time. Appendix A.1 describes the quality control measures applied to LLM-generated stories.
@@ 4 Human Study
Participant Recruiting. We recruited undergraduate students in English and Psychology from a US public university using targeted fliers and emails. We hypothesized that participants with some aptitude for literary and psychological analysis would provide more valuable insights than a random cross-section of the population. From 47 applications, we selected the 5 most promising candidates based on their interests and previous experience in narrative and psychological analysis. Our goal was to engage informed laypeople with relevant backgrounds, bridging the gap between typical Amazon Mechanical Turkers and expert professionals.
Evaluation Protocol. We held an initial meeting to introduce the PDS and outline the annotation task. We provided a tutorial annotation session, followed up with short questions designed to help participants calibrate their understanding of each component of the PDS. We show our annotation instructions in Appendix A.5. After confirming task comprehension, participants were instructed to complete the annotations remotely within a week. To prevent annotator fatigue and promote careful annotation, stories were divided into batches of 20. The evaluation criteria required participants to (1) read the prompts and stories thoroughly; (2) rate the five components of psychological depth on a Likert scale from 1 to 5; (3) assess the likelihood of authorship on a Likert scale from 1 (LLM) to 5 (human); (4) provide explanations for ratings (optional).
@@ 5 Results
Acquiring human annotations is often costly and time-consuming. Recent work has demonstrated the potential of leveraging LLMs for automated evaluation of text summarization Liu et al. (2023) and creative generation Rajani et al. (2023). Correspondingly, we designed an automated evaluation procedure to study the degree to which four contemporary LLMs of various sizes and inference costs – Llama-3-8B, Llama-3-70B, GPT-3.5 (gpt-3.5-turbo-0125), and GPT-4o (gpt-4o-2024-05-13) – can assess psychological depth in a zero-shot fashion. For each story, we prompted LLMs with instructions similar to those provided to human participants but additionally required explanations that contextualize each numerical rating.
@@
Overall, the average percentage increase due to Mixture-of-Personas prompting across all models is approximately . These results demonstrate that diverse LLM opinions can more accurately reflect the multifaceted nature of human judgment, proving potential for more nuanced and human-like assessments by AI systems.
@@
To compare human- and LLM-written stories, we aggregated participant ratings by author and present the means and standard deviations in Table 3. We also computed statistical significances via pairwise t-tests between each combination of authors. Full results for that analysis can be found in Appendix A.8.
@@
Beyond the five PDS components, participants were tasked with estimating authorship sources. Stories penned by humans averaged a rating of , compared to for the most "human-like" of LLMs, GPT-4. This discrepancy underscores the exceptional depth scores, particularly in light of the slightly higher suspicion of LLM authorship.
@@ 7.1 Creative Generation by LLMs
Other studies have concentrated moreso on fully automating the creative writing process. For example, the Weaver project Wang et al. (2024b) launched a series of LLMs that were meticulously pre-trained and fine-tuned with a focus on creative writing. With a maximum size of 70B parameters, Weaver Ultra was shown to outperform larger generalist LLMs when evaluated for style, relevance, fluency, and creativity. In addition to innovations in training, content-planning Yao et al. (2018); Goldfarb-Tarrant et al. (2019) and novel prompting strategies Yang et al. (2022) have been used to improve the factuality and coherence of creative generations, which we view as two necessary prerequisites for components of psychological depth like authenticity and narrative complexity.
7.2 Evaluating Creative Writing
Numerous studies have previously established methodologies for evaluating creative works, employing strategies ranging from the unstructured feedback of human experts Baer and Kaufman (2019), the application of specific evaluation rubrics Vaezi and Rezaei (2019), to the utilization of LLMs as autonomous critics Ke et al. (2023). Our research shares the closest methodological resemblance with a recent investigation by Chakrabarty et al. (2023). In that study, the authors proposed the Torrance Test of Creative Writing (TTCW) as a rubric to evaluate short stories for fluency, flexibility, originality, and elaboration. This assessment was applied to a corpus comprising 12 narratives authored by professional writers and 36 narratives produced by popular LLMs OpenAI et al. (2023); Anthropic (2023). Their findings showed that narratives authored by humans were 3-10 more creative than those generated by LLMs. Our research, however, arrives at a notably divergent conclusion regarding the creative capabilities and depth of LLMs, which we believe may be attributed to several methodological variances. Primarily, our analysis concentrates on stories of a considerably reduced length (450 vs. 1400 words), a decision influenced by the observed challenges LLMs face in producing lengthy texts seamlessly in a single iteration Yang et al. (2022). Additionally, the approach of iteratively regenerating stories to meet a specific word count could potentially detract from their overall quality. Moreover, we posit that the enhanced depth observed in our study may be due to more complex premises and prompting strategies, providing a richer framework for creativity compared to the simpler, single-sentence premises utilized in the aforementioned study.

### v2

@@ Abstract
Evaluations of creative stories generated by large language models (LLMs) often focus on objective properties of the text, such as its style, coherence, and diversity. While these metrics are indispensable, they do not speak to a story’s subjective, psychological impact from a reader’s perspective. We introduce the Psychological Depth Scale (PDS), a novel framework rooted in literary theory that measures an LLM’s ability to produce authentic and narratively complex stories that provoke emotion, empathy, and engagement. We empirically validate our framework by showing that humans can consistently evaluate stories based on PDS ( Krippendorff’s alpha). We also explore techniques for automating the PDS to easily scale future analyses. GPT-4o, combined with a novel Mixture-of-Personas (MoP) prompting strategy, achieves an average Spearman correlation of with human judgment while Llama-3-70B with constrained decoding scores as high as for empathy. Finally, we compared the depth of stories authored by both humans and LLMs. Surprisingly, GPT-4 stories either surpassed or were statistically indistinguishable from highly-rated human-written stories sourced from Reddit. By shifting the focus from text to reader, the Psychological Depth Scale is a validated, automated, and systematic means of measuring the capacity of LLMs to connect with humans through the stories they tell.
@@
Fabrice Harel-Canada Hanyu Zhou Sreya Muppalla Zeynep Yildiz Miryung Kim Amit Sahai‡ Nanyun Peng‡ University of California, Los Angeles fabricehc@cs.ucla.edu
Our code and data is available at https://github.com/PlusLabNLP/psychdepth.
@@ 1 Introduction
Stories play a crucial role in our understanding of ourselves and the world around us Langer (1942); Kaniss (1991). As large language models (LLMs) are increasingly deployed in narrative design and creation, their growing impact on how stories are told calls for a deeper understanding of their narrative power. Current evaluations of LLM-generated stories often focus on objective properties of the text such as discourse structure Liu et al. (2024), fluency Gómez-Rodríguez and Williams (2023), style Wang et al. (2024), creativity Chakrabarty et al. (2024), diversity Gehrmann et al. (2021), toxicity and bias Wang et al. (2023). However, it is crucial to extend these evaluations to accommodate the subjective, psychological impact stories have on readers. While some studies have recently explored aspects like empathy Ma et al. (2020); Sabour et al. (2022) and engagement Ghazarian et al. (2020); Xu et al. (2022), they do not fully capture the multifaceted and interconnected nature of the reading experience.
@@
To empirically validate the PDS, we conducted a study involving non-expert humans, who are increasingly engaging with creative content generated by LLMs. We enlisted five undergraduate students from UCLA to analyze a dataset of 97 stories authored by humans and five contemporary LLMs. The raters provided psychological depth ratings and predictions on human or machine authorship, with detailed justifications. Our approach, illustrated in Figure 2, addresses three critical research questions to explore the intersection of LLMs and the psychological nuances of creative writing.
@@
To what extent can psychological depth be measured automatically? Leveraging our novel Mixture-of-Personas prompting strategy, GPT-4o achieved a Spearman correlation of with human judgment while Llama-3-70B with constrained decoding attains correlations as high as for empathy and for narrative complexity. These results highlight that while no single LLM excels at predicting all components of psychological depth, a strategic combination of different LLMs shows significant promise for automating PDS analyses.
@@
Overall, our findings validate the Psychological Depth Scale as an effective, automated, and systematic means of measuring how well LLMs connect with humans through storytelling. Remarkably, our results reveal that GPT-4 already matches or exceeds the quality of respected stories from Reddit, with of readers believing GPT-4’s stories to be human-written.
@@ 2 The Psychological Depth Scale
The metrics for PDS are derived from an extensive literature review within cognitive psychology, media studies, and narrative analysis. We conducted a comprehensive search on several databases, including Google Scholar, PubMed, and JSTOR. Our search terms included “psychological depth in literary texts” and “cognitive narrative analysis”. The initial search results were screened for discussion and application of evaluative criteria for narrative quality and reader responses. Overall, we surveyed 95 peer-reviewed articles and books in the final review and extracted 143 candidate components from the included works. We employed thematic analysis to group these candidates under five metrics: (1) emotion provocation, (2) empathy, (3) engagement, (4) authenticity, and (5) narrative complexity. Below we analyze these metrics and discuss their impact. We note that while factors contributing to each metric may be complex, measuring their narrative achievement is relatively simple. As PDS is a reader-centered assessment tool, each metric is such that readers will know when a story achieves it. Our surveys reflect this fact.
Emotion Provocation (PROV) measures the narrative’s ability to elicit strong emotional responses, positive or negative. Recent fMRI research shows that congruent (i.e. positive valence - high intensity, negative valence - low intensity) textual emotive expressions are more cognitively effective than conflicting (i.e. positive valence - low intensity, negative valence - high intensity) ones Citron et al. (2014); Megalakaki et al. (2019). The disparity between congruent and conflicting narrative emotions may be a factor in "compassion fatigue" Maier (2015); Kinnick et al. (1996) and "psychic numbing" Maier et al. (2016); Slovic (2007) exhibited by readers towards news about mass tragedies. Enhancing a story’s emotional impact is therefore more complex than increasing the amount and intensity of emotion-laden content as this may elicit weaker emotional responses. Given its impact and complexity, we contend that emotional provocation is an achievement of psychologically deep stories and is thereby a crucial metric in narrative quality assessment.
Theories of emotions range across neuroscience Clark-Polner et al. (2016); Mendes (2016), cognitive psychology Ortony et al. (2022); Lazarus (1991), and philosophy Sartre (1971); Nussbaum (2004). Contemporary approaches describe emotions as mental states marked by valence (positivity or negativity) and arousal (level of intensity) Barrett et al. (2016). Narrative evocation of emotions foster attention and interest, and evoked emotions may persist or recur after reading Mar et al. (2011).
@@
Narrative Complexity (NCOM) refers to the presence of rich and intricate storylines and character development, especially those that engender puzzled intrigue from the reader Somasundaran et al. (2018). Narrative complexity is often achieved through creative techniques such as nonlinear narration, plot twists, and double perspectivation Kiss and Willemsen (2017). Complex narratives present immersive puzzles that motivate readers to undertake mental restructuring and retroactive revision Kiss and Willemsen (2017); Cutting (2019). Exerting cognitive effort facilitates reader attention and interest Empson (1947); Steiner (1978), and can result in a more enjoyable reading experience Zunshine (2006). However, recent studies show that LLM-generated stories often lack this kind of narrative depth, though improvements can be made by incorporating key discourse elements like suspense and diversity Tian et al. (2024). Additionally, planning and contextualizing narratives remain challenging for LLMs when compared to human authors Spangher et al. (2024).
Narrative complexity is also not narrational complexity: simple stories can be obscurely narrated, while complex stories can be free of complex language Cutting (2019). Furthermore, the extent to which narrative complexity is intersubjective depends on situational similarities and contextual cues Tikka and Kaipainen (2017). Thus, narrative complexity is a dynamic and reciprocal measure that goes beyond stylistic and structural choices.
@@ 3.1 Human Stories
We collected human-authored stories from Reddit’s r/WritingPrompts,111https://www.reddit.com/r/WritingPrompts a popular online community with over 18 million users. This forum was chosen for its accessibility, thematic diversity, and the structured nature of its writing prompts. Aspiring writers respond to these prompts (called "premises") with their stories, which readers can upvote or downvote. While these stories may not represent the pinnacle of human writing, Reddit’s voting system allows us to approximate where LLM performance stands relative to variances in human writing quality. We categorized the stories based on their ranking position: Human-Advanced for top-voted stories, Human-Intermediate for medium-voted stories, and Human-Novice for low-voted stories, with average upvotes of 1434, 263, and 9, respectively. To ensure that quality was the primary factor influencing votes, we only selected stories that were posted within 24 hours of one another.
@@ 3.2 LLM Stories
When generating stories with LLMs, we employed a multifaceted approach involving five models, two prompting strategies, and three sampled generations. We intentionally restricted the model architecture to examine the impact of model size on psychological depth. In particular, we chose the Llama-2 family Touvron et al. (2023) with a variety of sizes (7B, 13B, and 70B). We added Vicuna-33B Chiang et al. (2023), which is fine-tuned based on Llama, to fill the size gap between 13B and 70B. We also included GPT-4 Achiam et al. (2023) as the highest-performing LLM at the time (November 2023). Appendix A.1 describes the quality control measures applied to LLM-generated stories.
@@ 4 Human Study
Participant Recruiting. In November 2023, we recruited undergraduate students from UCLA’s English and Psychology departments using targeted fliers and emails. We hypothesized that participants with some aptitude for literary and psychological analysis would provide more valuable insights than a random cross-section of the population. From 47 applications, we selected the 5 most promising candidates based on their interests and previous experience in narrative and psychological analysis. Our goal was to engage informed laypeople with relevant backgrounds, bridging the gap between typical Amazon Mechanical Turkers and expert professionals.
Evaluation Protocol. In December 2023, we held an initial meeting to introduce the PDS and outline the annotation task. We provided a tutorial annotation session, followed up with short questions designed to help participants calibrate their understanding of each component of the PDS. We show our annotation instructions in Appendix A.5. After confirming task comprehension, participants were instructed to complete the annotations independently and remotely within seven days. To prevent annotator fatigue and promote careful annotation, stories were divided into batches of 20. The evaluation criteria required participants to (1) read the prompts and stories thoroughly; (2) rate the five components of psychological depth on a Likert scale from 1 to 5; (3) assess the likelihood of authorship on a Likert scale from 1 (LLM) to 5 (human); (4) provide explanations for ratings (optional).
@@ 5 Results
Acquiring human annotations is often costly and time-consuming. Recent work has demonstrated the potential of leveraging LLMs for automated evaluation of text summarization Liu et al. (2023) and creative generation Rajani et al. (2023). Correspondingly, we designed an automated evaluation procedure to study the degree to which four contemporary LLMs of various sizes and inference costs – Llama-3-8B, Llama-3-70B, GPT-3.5 (gpt-3.5-turbo-0125), and GPT-4o (gpt-4o-2024-05-13)222Llama-3 and GPT-4o were released after our human study ended, so we could only include them in the evaluation automation for RQ2. – can assess psychological depth in a zero-shot fashion. For each story, we prompted LLMs with instructions similar to those provided to human participants but additionally required explanations that contextualized each numerical rating.
To streamline interactions with these models, we used langchain Chase (2022) for querying OpenAI’s GPT-series and the guidance framework Lundberg et al. (2022) to support constrained decoding of the locally-hosted Llama-3 models. guidance provides enhanced LLM steerability that we used to guarantee parsable responses, thereby freeing the model to focus entirely on the depth annotation task without splitting attention on output formatting. We theorized that this separation of concerns would enable competitive performance relative to the proprietary GPT series, which does not expose underlying token probabilities and is, therefore, less well-suited for constraint-guided generation.
@@
Overall, the average percentage increase due to Mixture-of-Personas prompting across all models is approximately . These results demonstrate that diverse LLM opinions can more accurately reflect the multifaceted nature of human judgment, proving the potential for more nuanced and human-like assessments by AI systems.
@@
To compare human- and LLM-written stories, we aggregated participant ratings by author and present the means and standard deviations in Table 3. We also computed statistical significances via pairwise t-tests between each combination of authors. Full results for that analysis are shown in Appendix A.8.
@@
Beyond the five PDS components, participants were tasked with estimating authorship sources. Stories penned by humans averaged a rating of , compared to for the most "human-like" of LLMs, GPT-4. Despite being perceived as slightly less human on average, GPT-4’s psychological depth scores were generally higher on average. On the other hand, human-authored stories were correctly perceived as more human, but that was sufficient for them to garner better PSD scores.
@@ 7 Related Work
7.1 Evaluating Creative Writing
Numerous studies have established methodologies for evaluating creative works, ranging from unstructured feedback by human experts Baer and Kaufman (2019) and the use of specific evaluation rubrics Vaezi and Rezaei (2019), to employing LLMs as autonomous critics Ke et al. (2024). The Consensual Assessment Technique (CAT) Amabile (1982), widely regarded as the gold standard for subjectively evaluating creative works Baer and McKool (2014); Carson (2019); Baer (2017), traditionally relies on expert judgments to ensure reliable evaluations. However, our PDS framework aims to capture more universal psychological reactions to written works, so we broaden CAT by including non-experts. Given that AI-generated content is consumed by a diverse audience, incorporating non-experts allows us to better characterize a wider range of psychological and emotional responses.
Methodologically, our research shares the closest resemblance with a recent investigation by Chakrabarty et al. (2024). In that study, the authors proposed the Torrance Test of Creative Writing (TTCW) as a rubric to evaluate short stories for fluency, flexibility, originality, and elaboration. This assessment was applied to a corpus comprising 12 narratives authored by professional writers and 36 narratives produced by popular LLMs Achiam et al. (2023); Anthropic (2023). Their findings showed that narratives authored by humans were 3-10 more creative than those generated by LLMs. Our research, however, arrives at a notably divergent conclusion regarding the creative capabilities and depth of LLMs, which we believe may be attributed to several methodological variances. Primarily, our analysis concentrates on stories of a considerably reduced length (450 vs. 1400 words), a decision influenced by the observed challenges LLMs face in producing lengthy texts seamlessly in a single iteration Yang et al. (2022). Additionally, the approach of iteratively regenerating stories to meet a specific word count could potentially detract from their overall quality. Moreover, we posit that the enhanced depth observed in our study may be due to more complex premises and prompting strategies, providing a richer framework for creativity compared to the simpler, single-sentence premises utilized in the aforementioned study.
@@ 7.2 Creative Generation by LLMs
Other studies have concentrated moreso on fully automating the creative writing process. For example, the Weaver project Wang et al. (2024) launched a series of LLMs that were meticulously pre-trained and fine-tuned with a focus on creative writing. With a maximum size of 70B parameters, Weaver Ultra was shown to outperform larger generalist LLMs when evaluated for style, relevance, fluency, and creativity. In addition to innovations in training, content-planning Yao et al. (2019); Goldfarb-Tarrant et al. (2019) and novel prompting strategies Yang et al. (2022) have been used to improve the factuality and coherence of creative generations, which we view as two necessary prerequisites for components of psychological depth like authenticity and narrative complexity.